# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks: The GATTACA Framework

### v1

@@ Abstract
Cellular reprogramming, the artificial transformation of one cell type into another, has been attracting increasing research attention due to its therapeutic potential for complex diseases. However, discovering reprogramming strategies through classical wet-lab experiments is hindered by lengthy time commitments and high costs.
@@
In this study, we explore the use of deep reinforcement learning (DRL) to control Boolean network models of complex biological systems, such as gene regulatory networks and signalling pathway networks. We formulate a novel control problem for Boolean network models under the asynchronous update mode in the context of cellular reprogramming. To facilitate scalability, we consider our previously introduced concept of a pseudo-attractor and we improve our procedure for effective identification of pseudo-attractor states. Finally, we devise a computational framework to solve the control problem. To leverage the structure of biological systems, we incorporate graph neural networks with graph convolutions into the artificial neural network approximator for the action-value function learned by the DRL agent. Experiments on a number of large real-world biological networks from literature demonstrate the scalability and effectiveness of our approach.
1 Introduction
@@
Complex diseases remain a major challenge for biomedical research due to the intricate interactions between genes forming gene regulatory networks (GRNs). This network exhibits complex dynamics, emerging from the structure of the network: the state of the genome evolves in time, finally reaching some stable gene expression profiles characterising cell types or, more generally, stable cellular functional states [11].
@@
In some cases, when a GRN is perturbed, it can reach an “unhealthy”, undesired state, that is otherwise inaccessible [5]. An artificial change of the cell fate – called cellular reprogramming aims to solve this issue by steering the GRN back to a “healthy” state. However, identifying effective reprogramming interventions through wet-lab experiments alone is prohibitively expensive and time-consuming. This motivates the development of in-silico methods. To this end, GRNs can be holistically modelled with the formalism of Boolean networks (BNs) originally introduced by S. Kauffman [14]. Its simplicity enables tractable simulations and analysis of non-linear biological dynamics while retaining essential, emergent system behaviour.
@@
To support cellular reprogramming, control strategies must be computed for large-scale BNs, taking into account both network structure and dynamics [9]. Existing state-of-the-art methods, such as CABEAN [30], provide exact algorithms for attractor-based control, but these approaches are limited in scalability due to the exponential growth of the state space and assumptions about network decomposability.
Deep reinforcement learning (DRL) is a promissing alternative for traditional methods for controlling GRNs. It was pioneered in [27] which focused on avoiding undesirable states in terms of steady state probabilities of PBNs. The authors treated gene expression samples as a time-series of experience tuples and used Q-Learning to approximate the optimal policy. To relax the temporal condition, the BOAFQI-Sarsa method was devised in [22]. A further development was mSFQI method proposed in [21], which once again was for control based on the probabilities of gene activity profiles.
The first method that attempted exact control of BNs was the eXtended Classifier System (XCS), which was rule-based reinforcement learning applied to the control of random BNs in [12], and subsequently to the 11-node yeast cell cycle network from [15] in [13].
Scalability is the main concern when devising new methods for control of BNs. All of the methods descirbed above could only control BNs of limited size, with the exact method stopping at 11 nodes. Since deep reinforcement learning (DRL) techniques have proven highly successful in decision problems characterised by large state-action spaces the idea of combining Q-learning with deep learning was conceived. The first ones to realise it were Papagiannis et al. in [23]. They proposed a control method based on deep Q-Learning, namely Double Deep Q-Network (DDQN) with Proritised Experience Replay (PER), initialy introduced in [8]. This method was successful in [24] to control a 20-node synthetic PBN and a 7-node real-world PBN model of melanoma. The same method was used by Acernese et al. in [1] for the control of synchronous PBNs with the goal of driving the networks from a particular state towards a more desirable one, and also in [2] to address the feedback stabilisation problem of 9-gene probabilistic Boolean control networks (PBCNs) modeling apoptosis with the aim of stabilising the system at a given equilibrium point. To tackle the control problem on large networks Moschoyiannis et al. in [19] applied DDQN to control a large synchronous melanoma PBN model of 200 nodes. To make their method scale they limited controll to flipping a single, specific gene, namely pirin, which had already been identified as relevant for the control of PBN models of melanoma in previous studies. Also, the target set of states consisted of exactly half of the entire state space, i.e. the objective was to reach a subset of network states where the WNT5A gene is OFF.
@@
In our previous work [17], we introduced pbn-STAC – a BDQ-based tool for control of asynchronous BNs and probabilistic Boolean networks (PBNs). Therein, we introduced a notions of a pseudo-attractor and a pseudo-attractor state, we devised a pseudo-attractor states identification procedure (PASIP), and we formulated a novel control problem, i.e. the Source-Target Attractor Control, that corresponds with the problem of identifying cellular reprogramming strategies. Our pbn-STAC framework was shown to be able to find optimal and sub-optimal control strategies between source and target attractors for BN and PBN models of sizes up to 33 nodes.
In this study, we improve on our previous work to be able to compute control strategies for very large BN models of biological networks. To this end, we introduce the Graph-based Attractor-Target Control Algorithm (GATTACA), a DRL framework designed to solve a novel target control problem for BN models of GRNs under the asynchronous update mode. As in [17], and unlike other previous approaches, we restrict control actions to be applied only in observable phenotypic cellular states, which are hypothesised to be represented by pseudo-attractors of a BN model. To identify these efficiently, we enhance our earlier pseudo-attractor states identification procedure to handle large and complex networks.
@@
Our method uses a Graph Neural Network (GNN)-based Q-learning agent (ATGC) that learns minimal control strategies to drive the system from a pseudo-attractor state to one aligned with a given target configuration, i.e. specified value for a subset of nodes. The structure of the GRN is directly encoded into the GNN, allowing the model to exploit topological relationships between genes during training.
@@
To assess the scalability potential, we evaluated GATTACA on several large-scale, real-world biological networks, demonstrating its scalability and effectiveness in cases where exact symbolic approaches, such as those implemented in CABEAN, fail. Our contributions represent a step forward in developing predictive, scalable and effective computational methods for cellular reprogramming scenarios.
@@
The paper is organised as follows. The preliminaries are presented in Section 2. Our novel target control problem is the context of cellular reprogramming in formulated in Section 3. Next, a DRL-based framework for controlling large-scale biological networks is developed in Section 4. The framework is evaluated on large models of real-world gene regulatory and signalling networks in accordance with a methodology presented in Section 5. The obtained evaluation results are discussed in Section 6. Finally, the the study is concluded in Section 7.
2 Preliminaries
2.1 Boolean networks
We formally define a Boolean network (BN) as a discrete dynamical system represented by a directed graph $G=(V,E)$, where $V$ denotes the set of nodes representing elements within the system, and $E$ represents the set of directed edges denoting interactions between these elements. Each node $v_{i}$ is associated with a Boolean function $f_{i}$, expressing the state transition rule of the node. These Boolean functions determine the state of each node at time $t+1$, based on the states of its input nodes at time $t$. Mathematically, $f_{i}:\{0,1\}^{k}\rightarrow\{0,1\}$, where $k$ denotes the number of input nodes for node $v_{i}$. The system dynamics evolve iteratively, with the states of the nodes being updated synchronously or asynchronously according to their respective Boolean functions.
@@ 2.1 Boolean networks
A state $x$ of a Boolean network $G=(V,E)$ is a binary vector $x=(x_{1},x_{2},...,x_{n})$, where $n$ is the number of nodes in the network $|V|$, and each $x_{i}$ represents the state of node $v_{i}$ at a particular time step.
@@
A state transition graph (STG) for a Boolean network $G=(V,E)$ is a directed graph that represents the possible state transitions of the network. The nodes of STG correspond to states of $G$, and there exists a directed edge from state $s$ to state $s^{\prime}$ in STG if and only if there exists a Boolean function in $G$ such that applying it to state $s$ results in state $s^{\prime}$. Formally, $STG=(S,\rightarrow)$, where $S$ is the set of all possible states of $G$, and $\rightarrow$ is the set of directed edges representing state transitions.
@@
An attractor of a BN is a bottom strongly connected component (BSCC) in the STG of the network. A fixed-point attractor and a multi-state attractor are BSCCs consisting of a single state or more than one state, respectively.
@@
In the case of the asynchronous update mode, the dynamics of a BN is non-deterministic. It can be modelled as a Markov chain. In this view, an attractor of the BN corresponds to an irreducible subset of states of the Markov chain.
2.2 Pseudo-attractors
@@
It is proven that obtaining the family of all attractors of a BN is NP-hard by itself [3]. This is especially prohibitive in the case of cellular reprogramming considerations, where ultimately large models need to be analysed. However, as we also argued in [17], from the practical point of view, in the context of cellular reprogramming only the frequently revisited states of an attractor are the relevant ones since they correspond to phenotypical cellular states that are observable in the lab. This makes these states ‘recognisable’ for the application of reprogramming interventions in practice in accordance with the control strategy obtained with our framework. With this in mind, we have introduced in [17] the notions of a pseudo-attractor and a pseudo-attractor state in the case of the more general framework of probabilistic Boolean networks (PBNs). Moreover, we proposed the Pseudo-Attractor States Identification Procedure (PASIP) that facilitates the determination pseudo-attractor states in large models. Since in this study our focus is on BN models under the asynchronous update mode, herein we provide a specific formulation of the definition of a pseudo-attractor and a pseudo-attractor state restricted to asynchronous BNs as follows.
Definition 2.1 (Pseudo-attractor).
@@
Let $A$ be an attractor of an asynchronous BN, i.e. an irreducible set of states of the Markov chain underlying the BN. Let $n:=|A|$ be the size of the attractor $A$ and let $\mathbb{P}_{A}$ be the unique stationary probability distribution on $A$. The pseudo-attractor associated with $A$ is the maximal subset $PA\subseteq A$ such that for all $s\in PA$ it holds that $\mathbb{P}_{A}(s)\geq\frac{1}{n}$. The states of a pseudo attractor are referred to as pseudo-attractor states.
@@
As argued in [17], the correctness of the definition is guaranteed by the fact that the state space of the underlying Markov chain of a BN is finite and that the Markov chain restricted to the attractor states is irreducible. It is a well known fact that all states of a finite and irreducible Markov chain are positive recurrent. In consequence, the attractor-restricted Markov chain has a unique stationary distribution.
2.3 Reinforcement learning
The Attractor-Target Control can be interpreted as a Markov Decision Process (MDP) and it can be solved using reinforcement learning (RL). RL solves sequential decision problems by optimizing a cumulative reward function.
The RL problem is meant to be a direct framing of the task of learning from interaction to achieve a goal. The decision-maker and learner is referred to as the agent. At each time step $t$, the agent interacts with the environment which comprises everything outside the agent by implementing a mapping $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ from states to probabilities of selecting an action $A_{t}$ in state $S_{t}$. This mapping is referred to as the agent’s policy, where $\pi(a\mid s)$ is the probability that $A_{t}=a$ given that $S_{t}=s$.
Within the MDP-based framework, the RL problem is formalised as the optimal control of MDPs. RL methods specify how the agent changes its policy as a result of its experiences with the goal of maximising the expected discounted return:

### v2

# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks – the GATTACA Framework

@@ Abstract
Cellular reprogramming, the artificial transformation of one cell type into another, has been attracting increasing research attention due to its therapeutic potential for complex diseases. However, identifying effective reprogramming strategies through classical wet-lab experiments is hindered by lengthy time commitments and high costs.
@@
In this study, we explore the use of deep reinforcement learning (DRL) to control Boolean network models of complex biological systems, such as gene regulatory and signalling pathway networks. We formulate a novel control problem for Boolean network models under the asynchronous update mode, specifically in the context of cellular reprogramming. To solve it, we devise GATTACA, a scalable computational framework.
To facilitate scalability of our framework, we consider previously introduced concept of a pseudo-attractor and improve the procedure for effective identification of pseudo-attractor states. We then incorporate graph neural networks with graph convolution operations into the artificial neural network approximator of the DRL agent’s action-value function. This allows us to leverage the available knowledge on the structure of a biological system and to indirectly, yet effectively, encode the system’s modelled dynamics into a latent representation.
Experiments on several large-scale, real-world biological networks from the literature demonstrate the scalability and effectiveness of our approach.
Index Terms:
Boolean Networks, Network Control, Deep Reinforcement Learning, Gene Regulatory Networks, Cellular Reprogramming
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Complex diseases remain a major challenge for biomedical research due to the intricate interactions between genes forming gene regulatory networks (GRNs). These networks exhibit complex dynamics emerging from the structure of the network: the state of the genome evolves in time, finally reaching some stable gene expression profiles characterising cell types or, more generally, stable cellular functional states [1].
@@
In some cases, when a GRN is perturbed, it can reach an “unhealthy” state, that is otherwise inaccessible [2]. An artificial change of the cell fate, referred to as cellular reprogramming, aims to solve this issue by steering the GRN back to a “healthy” state. However, identifying effective reprogramming targets and intervention strategies through wet-lab experiments alone is prohibitively expensive and time-consuming. This motivates the development of in-silico methods.
Capturing emergent properties of GRN dynamics requires the consideration of the system as a whole. To this end, GRNs can be holistically modelled with the formalism of Boolean networks (BNs) originally introduced by Kauffman in [3]. Its simplicity enables tractable simulations and analysis of non-linear biological dynamics while retaining essential, emergent complex system behaviour. Cellular reprogramming can be formulated as a control problem within the framework of BNs.
@@
To support cellular reprogramming, control strategies must be computed for large-scale BNs, taking into account both the structure and dynamics of the network [4]. Scalability is the main concern when devising new computational methods for control of asynchronous BNs. Existing state-of-the-art methods and tools, such as CABEAN [5], provide exact algorithms for attractor-based control. Nevertheless, they are limited in scalability because of the infamous “curse of dimensionality” problem, i.e. the size of the state space grows exponentially with the number of genes considered, which is handled by additional assumptions about network decomposability.
@@
Deep reinforcement learning (DRL) has proven highly successful in decision problems characterised by very large state-action spaces [6]. Therefore, it seems a promising alternative to traditional methods for controlling GRNs. In [7], pbn-STAC – a DRL-based tool for control of asynchronous BNs and probabilistic Boolean networks (PBNs) was introduced, which was a result of exploration into the potential of DRL for developing scalable control methods for both BN and PBN models of GRNs in the context of cellular reprogramming. Therein, notions of a pseudo-attractor and a pseudo-attractor state were introduced, a pseudo-attractor states identification procedure (PASIP) was devised, and a novel control problem, i.e. the Source-Target Attractor Control, that corresponds with the problem of identifying cellular reprogramming strategies was formulated. The pbn-STAC framework was shown to be able to find optimal and sub-optimal control strategies between source and target attractors for BN and PBN models of sizes up to 33 nodes.
In this study, we build on the initial idea of [7] to devise a general and scalable approach capable of handling very large BN models of biological networks. To this end, we introduce the Graph-based Attractor-Target Control Algorithm (GATTACA) computational framework to solve a novel, general target-control problem for BN models of biological networks under the asynchronous update mode. Our new control problem is tailored for cellular reprogramming in the sense that control actions are allowed to take place only in attractor states as they correspond to observable phenotypic cellular states understood as discrete, mutually exclusive, stable, and observable cell phenotypes or cell fates that can undergo transitions from one to another under the influence of regulatory inputs. This approach addresses the constraints of wet-lab biology, where it is very difficult to experimentally determine a transient state of a cell, without destroying it, for the application of a control intervention in a real-world scenario. In contrast, stable cell phenotypes, which correspond to the attractor states of a dynamical system [8], are expected to be more easily, optically observable realisations of complex interactions between the genome, epigenome, and local environment. To reflect these constraints, we restrict control actions to be applied only in observable phenotypic cellular states, which are represented by attractors of a BN model.
@@
The core of our GATTACA framework is a Graph Neural Network (GNN)-based Q-learning agent which learns minimal control strategies that drive the system from a given source attractor state to one aligned with a given target configuration, i.e. an attractor with user-specified values for a subset of nodes. The structure of the biological network is encoded directly into the GNN, allowing the model to exploit topological relationships between the network components (nodes) during training. Furthermore, to facilitate effective identification of attractors in large BN models of complex networks, we consider the notion of pseudo-attractor states introduced in [7] and we enhance the pseudo-attractor states identification procedure proposed therein.
@@
To assess the scalability potential, we evaluate the GATTACA framework on several large-scale, real-world biological networks, demonstrating its scalability and effectiveness in cases where exact symbolic approaches, such as those implemented in CABEAN, fail. Our contributions result in a scalable and effective computational method for cellular reprogramming scenarios.
@@
The paper is organised as follows. Related works are discussed in Section II. The preliminaries are presented in Section III. Our novel target control problem in the context of cellular reprogramming is formulated in Section IV. Next, a DRL-based framework for controlling large-scale biological networks is developed in Section V. The framework is evaluated on large models of real-world gene regulatory and signalling networks in accordance with a methodology presented in Section VI. The obtained evaluation results are discussed in Section VII. Finally, the study is concluded in Section VIII.
II Related works
The existing structure- and dynamics-based state-of-the-art computational techniques are limited to small and mid-size networks due to the infamous state space explosion problem. They usually require the systems to facilitate some kind of decomposition. However, this is often too restrictive for cellular reprogramming considerations.
The application of reinforcement learning to control GRNs was pioneered in [9]. The study focused on avoiding undesirable states in terms of steady-state probabilities of PBNs. The authors treated time-series gene expression samples as a sequence of experience tuples and used a batch version of Q-Learning to approximate the optimal policy. Later, to relax the temporal condition, the BOAFQI-Sarsa method was devised in [10]. A further development was mSFQI method proposed in [11] for control based on the probabilities of gene activity profiles.
Rule-based reinforcement learning in the form of an eXtended Classifier System (XCS) was devised in [12] for the control of random BNs. Subsequently, it was applied in [13] to a 11-node yeast cell cycle network.
Scalability is the main concern when devising new methods for control of BNs. This issue can be addressed by devising new methods based on deep reinforcement learning (DRL) techniques, which have proven highly successful in decision problems characterized by large state-action spaces, see, e.g. [6]. To this end, the idea of combining Q-learning with deep learning was conceived. The first to apply this idea in the context of BN control were Papagiannis et al. in [14], where a control method for BNs based on deep Q-Learning with experience reply, namely Double Deep Q-Network (DDQN) with Proritised Experience Replay (PER), was proposed. It was applied in [15] to control a 20-node synthetic PBN and a 7-node real-world PBN model of melanoma. The same approach was used by Acernese et al. in [16] for the control of synchronous PBNs with the goal of driving the networks from a particular state towards a more desirable one. Next, in [17], the authors proposed a Q-learning based method to address the feedback stabilisation problem of probabilistic Boolean control networks (PBCNs) with the aim of stabilising the system at a given equilibrium point and applied it to a small 9-node model of apoptosis.
To address the control problem on large networks, Moschoyiannis et al. in [18] applied DDQN with PER to control a large synchronous melanoma PBN model of 200 nodes. They limited the control actions to flipping a single specific gene, namely pirin, which had already been identified as relevant for the control of melanoma PBN models in previous studies. Furthermore, the target set of states consisted of exactly half of the entire state space, that is, the objective was to reach a subset of network states where the WNT5A gene is OFF.
Finally, Mizera et al. in [7] leveraged DRL to develop pbn-STAC, a computational framework that solves the source-target attractor control problem of driving the network dynamics from a given source to a specified target attractor of a BN or PBN model under the asynchronous update mode. The control problem is tailored for cellular reprogramming in the sense that control actions are allowed to take place only in intermediate attractor states to reflect the constraints of wet-lab biology. The approach was successful in controlling BN and PBN models of up to 33-nodes.
III Preliminaries
III-A Boolean networks
We formally define a Boolean network $B$ as a discrete dynamical system represented by a pair $B=(V,\mathcal{F})$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is a set of binary-valued nodes (a.k.a genes) and =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ is a set of predictor functions. Each predictor function $f_{i}\in\mathcal{F}$ is a Boolean function that encodes the state transition rule of node $v_{i}$. It is defined with respect to a subset of $V$ denoted $\textrm{Pa}(v_{i})\subseteq V$ and referred to as the parent nodes (regulators) of $v_{i}$ in $B$. The predictor function $f_{i}$ determines the state of $v_{i}$ at a discrete time point $t+1$ based on the states of the parent nodes at time point $t$.
structure graph of a BN is a directed graph $G=(V,E)$, where $V$ is the set of nodes of the BN representing components of the system and $E$ is a set of edges defined as $E=\{(v,u)\in V\times V\mid v\in\textrm{Pa}(u)\}$. The structure graph represents the individual direct regulatory relationships between the components of the system encoded by the respective Boolean functions.
The system dynamics evolves iteratively, with the states of the nodes being updated in accordance with their respective predictor functions. Two update modes are usually considered: synchronous, where all nodes are updated at once, and asynchronous, where one randomly chosen node is updated at a time. As in [7], also in this study we focus on the asynchronous mode, which is considered more appropriate for the modelling of biological networks [19, 20], yet more challenging from a computational point of view.
@@ III-A Boolean networks
state $x$ of a Boolean network $B=(V,\mathcal{F})$ is a binary vector $x=(x_{1},x_{2},\ldots,x_{n})$, where $n=|V|$ is the number of nodes in the network and each $x_{i}\in\{0,1\}$ represents the state of node $v_{i}$.
@@
An associated state transition graph (STG) of $B$ under the asynchronous update mode is a directed graph that represents the possible state transitions of $B$. Nodes of the STG correspond to states of $B$, and there exists a directed edge from state $s$ to state $s^{\prime}$ in the STG if and only if there exists a predictor function in $\mathcal{F}$ such that applying it to state $s$ results in state $s^{\prime}$. Formally, $\textrm{STG}(B)=(S,\rightarrow)$, where $S=\{0,1\}^{|V|}$ is the set of all possible states of $B$, and $\rightarrow$ is the set of directed edges representing state transitions, i.e. =$\rightarrow=\{(s,s^{\prime})\in S\times S\mid\exists f_{i}\in\mathcal{F}\ f_{i}(s)=s^{\prime}\}$.
@@
An attractor of a BN is a bottom strongly connected component (BSCC) in the STG of the network. A fixed-point attractor and a multi-state attractor are BSCCs consisting of a single state and more than one state, respectively.
@@
In the case of the asynchronous update mode, the dynamics of a BN is non-deterministic, i.e. $\rightarrow$ is a relation rather than a function. It can be modelled as a Markov chain. In this view, an attractor of the BN corresponds to an irreducible subset of states of the Markov chain.
For some examples of the introduced concepts related to BNs, we refer to Appendix -A.
III-B Pseudo-attractors
@@
It was shown that obtaining the family of all attractors of a BN is NP-hard by itself [21]. This is especially prohibitive in the case of cellular reprogramming considerations, where ultimately large models need to be analysed. However, from a practical point of view in the context of cellular reprogramming, only the frequently revisited states of an attractor are the relevant ones since they correspond to phenotypical cellular states that are observable in the lab. This makes these states ‘recognisable’ for the application of reprogramming interventions in practice in accordance with the control strategy obtained with our framework. With this in mind, the notions of a pseudo-attractor and a pseudo-attractor state were introduced in [7]. Since in this study our focus is solely on BN models under the asynchronous update mode, herein we provide a variant of the definition explicitly referring to asynchronous BNs.
Definition III.1 (Pseudo-attractor).
@@
Let $\mathcal{A}$ be an attractor of an asynchronous BN, i.e. an irreducible set of states of the Markov chain underlying the BN. Let $n:=|\mathcal{A}|$ be the size of the attractor $\mathcal{A}$ and let $\mathbb{P}_{\mathcal{A}}$ be the unique stationary probability distribution on $\mathcal{A}$. The pseudo-attractor associated with $\mathcal{A}$ is the maximal subset $\mathcal{PA}\subseteq\mathcal{A}$ such that for all $s\in\mathcal{PA}$ it holds that $\mathbb{P}_{\mathcal{A}}(s)\geq\frac{1}{n}$. The states of a pseudo attractor are referred to as pseudo-attractor states.
@@
The correctness of the definition is guaranteed by the fact that the state space of the underlying Markov chain of an asynchronous BN is finite and that the Markov chain restricted to the attractor states is irreducible. It is a well known fact that all states of a finite and irreducible Markov chain are positive recurrent. In consequence, the attractor-restricted Markov chain has a unique stationary distribution.
In [7] it was proved that for any BN attractor there exists a non-empty pseudo-attractor as stated by the following two theorems.
Theorem III.2.
Let $\mathcal{A}$ be an attractor of a BN. Then there exists a pseudo-attractor $\mathcal{PA}\subseteq\mathcal{A}$ such that $|\mathcal{PA}|\geq 1$.
Theorem III.3.
Let $\mathcal{A}$ be an attractor of a BN such that the unique stationary distribution of the underlying Markov chain of the BN restricted to $\mathcal{A}$ is uniform. Then, for the pseudo-attractor $\mathcal{PA}$ associated with $\mathcal{A}$ it holds that $\mathcal{PA}=\mathcal{A}$, i.e. the pseudo-attractor and the attractor are equal.
Moreover, we proposed the pseudo-attractor state identification procedure (PASIP) that facilitates the determination of PA states in large models. In this study, we identify a condition which may result in the blockage of the original PASIP. We provide a solution to this issue and present here the improved version of PASIP. The details of the method are presented and discussed in Section V-A.
III-C Reinforcement learning
A control problem for BNs can be viewed as a Markov Decision Process (MDP) and can be solved using reinforcement learning (RL). RL solves sequential decision problems by optimizing a cumulative reward function.
The RL problem is meant to be a direct framing of the task of learning from interactions to achieve a goal. The decision-maker, or learner, is referred to as an agent. At each discrete time step $t$, the agent interacts with an environment, which encompasses everything external to the agent. The agent’s behaviour is defined by a policy, given by a mapping $\pi:\mathcal{S}\times\mathcal{AC}\rightarrow[0,1]$ defined for all observable states in the environment $\mathcal{S}$ and all possible actions in action space $\mathcal{AC}$, where $\pi(a\mid s)$ is the probability of selecting action $a\in\mathcal{AC}$ while being in state $s\in\mathcal{S}$. This policy guides the agent’s decision-making process throughout its interaction with the environment.
Within the MDP framework, the RL problem is formalised as the optimal control of an MDP. RL methods specify how the agent changes its policy as a result of its experiences with the goal of maximising the expected discounted return: